OpenStudy (anonymous):

What is the definition of refelxive, symmetric,transitive,associative,commutative and distributive properties. Examples would be very helpful.

OpenStudy (blacksteel):

The reflexive property: a=a For example, 3 = 3 ----- The symmetric property: If a=b, then b=a For example, x = 3 implies that 3 = x ----- The transitive property: If a=b and b=c, then a=c For example, if x = 3 and 3 = y, then x=y ----- The associative property: (a + b) + c = a + (b + c) and (a*b)*c = a*(b*c) For example: (1+2)+3 = 3 + 3 = 6 = 1+5 = 1+(2+3) (2*3)*4 = 6*4 = 24 = 2*12=2*(3*4) ----- The commutative property: a + b = b+ a and a*b = b*a For example: 3+4 = 7 = 4+3 2*3 = 6 = 3*2 ----- The distributive property: a*(b+c) = a*b + a*c For example: 3*(4+5) = 3*9 = 27 = 12 + 15 = 3*4 + 3*5 ----- Hope this helps! If you have any questions, feel free to ask

OpenStudy (anonymous):

Thanks a bunch!